OpenStudy (anonymous):

Help with Trig!!-- 2 tan x +2=0....... solve for all degree solutions and x = 0 degrees < x< 360 degrees

OpenStudy (anonymous):

135/2 degs and 315/2 degs.

OpenStudy (anonymous):

maybe there are more answer.s..

OpenStudy (anonymous):

oops no. sry im wrong

OpenStudy (anonymous):

can you please show me how to solve this step by step? please

OpenStudy (anonymous):

im wrong

OpenStudy (anonymous):

i moved 2 over and divided by 2 to get tax=-1

OpenStudy (anonymous):

*tanx =-1

OpenStudy (anonymous):

and thats true when x = 135/2 degs and 315/2 degs.

OpenStudy (anonymous):

no no corss that, its true when 315 degs ad 135 degs..

OpenStudy (anonymous):

OHH WAIT

OpenStudy (anonymous):

oh so the answer is 135 and 315 degrees

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

only for TANX. you have 2tanx+2

OpenStudy (anonymous):

Tan is -1 at 180 - 45 degrees = 135 degrees and Tan is -1 at 360 - 45 degrees = 315 degrees

OpenStudy (anonymous):

no. plug into equation and it wont work

OpenStudy (anonymous):

oh i just looked in the back of my textbook... it is

OpenStudy (anonymous):

it only shows answers not steps.. i appreciate it

OpenStudy (anonymous):

im confused.. which one has the multiple angles ? like is it tan(x/2)

OpenStudy (anonymous):

ok nvm yeah

OpenStudy (anonymous):

First replace tan x with something else to make it easier to do. Then solve for that new variable. \[\text{Let tan x} = g \implies 2(tan\ x)+2 = 2g+2 \] Then solve for g. Then replace g with tan x again and take the inverse tan to find the angle between \(\frac{-\pi}{2}\) to \(\frac{\pi}{2}\) that satisfies that result. Then you can add or subtract additional \(\pi\) values to find more angles.

OpenStudy (anonymous):

i understand everything your telling me, howeverr.. im confused on why you subtracted 45 degrees

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

nvm.. i understand, thank you very much